OpenStudy (anonymous):

∫[∞on top and 1 on the bottom](Inx)/(√(x))dx

myininaya (myininaya):

\[\int\limits_{1}^{\infty}\frac{lnx}{\sqrt{x}}dx\]

OpenStudy (anonymous):

use integration by parts is it?

OpenStudy (anonymous):

no substitute lnx =t nd solve

OpenStudy (anonymous):

and for e denominator use substitution?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

the answer to your ques. is 4

OpenStudy (anonymous):

using l'hopital's rule?

OpenStudy (anonymous):

no substitution ....

myininaya (myininaya):

\[u=lnx, \frac{du}{dx}=\frac{1}{x}, du=\frac{1}{x} dx\] \[u=lnx, e^u=x, e^\frac{u}{2}=x^\frac{1}{2}\] \[\int\limits_{}^{}\frac{lnx}{x}\sqrt{x}dx=\int\limits_{}^{}u*e^\frac{u}{2}du=2e^\frac{u}{2}*u-\int\limits_{}^{}2e^\frac{u}{2}du=2e^\frac{u}{2}*u-4e^\frac{u}{2}+C\] \[=2e^\frac{lnx}{2}*lnx-4e^\frac{lnx}{2}+C=2e^{lnx^\frac{1}{2}}*lnx-4e^{lnx^\frac{1}{2}}+C\] \[=2x^\frac{1}{2}*lnx-4x^\frac{1}{2}+C\]

myininaya (myininaya):

i did substitution+integration by parts

myininaya (myininaya):

now we need to go back in deal with our limits

OpenStudy (anonymous):

yes,i got to this step. then sub. the values?

OpenStudy (anonymous):

yes dat was wat i was sayng n m too lazy to type all this in equations.. thx myininaya

myininaya (myininaya):

lol

myininaya (myininaya):

ok can you finish it mathstina?

myininaya (myininaya):

im gonna brush my teeth and i will be back

OpenStudy (anonymous):

In ∞= 0 right

OpenStudy (anonymous):

?

OpenStudy (cruffo):

no, log grows with out bound. I don't think the integral converges.

OpenStudy (anonymous):

i meant sub ∞ for that qn

OpenStudy (anonymous):

mathstina... wen u substitute x=inf int the integral wat u get is zero .. inf-inf \[ 2x ^{0.5}lnx-4x^{0.5}\]

OpenStudy (anonymous):

singh,tks

OpenStudy (cruffo):

\[\infty - \infty\] is indeterminant. \[2\lim_{x \to \infty}\; \sqrt{x}\;(\ln(x)-2) = \infty\]

OpenStudy (anonymous):

@cruffo... X is not approaching infinity x is infinity @mathstina Happy to help :)

OpenStudy (cruffo):

whatever....

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

thnks for everyone's help.

myininaya (myininaya):

the integral diverges

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

shall i ask 2qns on partial fractions? myininaya is it time for u to sleep?

myininaya (myininaya):

we have \[\lim_{b \rightarrow \infty}[(2b^\frac{1}{2}*lnb-4*b^\frac{1}{2})]-[2*1^\frac{1}{2}*\ln(1)-4*1^\frac{1}{2}]\] \[\lim_{b \rightarrow \infty}(2\sqrt{b}*lnb-4\sqrt{b}+4)=4+\lim_{b \rightarrow \infty}2\sqrt{b}*(lnb-2)\]

myininaya (myininaya):

\[=\infty\]

myininaya (myininaya):

ok post a new thread because this is getting laggy

OpenStudy (anonymous):

i thought the final ans is 4.

myininaya (myininaya):

let me look at it again

OpenStudy (anonymous):

am i right? is the ans infinity

OpenStudy (anonymous):

the answer is four guys......

OpenStudy (anonymous):

i checked the answer with my calculator its four :/

OpenStudy (anonymous):

∞+4=4?

OpenStudy (cruffo):

What did you type into the calculator?

myininaya (myininaya):

i get it diverges with my calculator

OpenStudy (anonymous):

or 0+4=4?

myininaya (myininaya):

infinity+4=infinity

OpenStudy (cruffo):

I agree with myininaya, the integral diverges.

myininaya (myininaya):

and i agree with cruffo

myininaya (myininaya):

lol

OpenStudy (cruffo):

:)

OpenStudy (anonymous):

singh,myninaya said the ans is infinity

myininaya (myininaya):

man, do you have maple?

myininaya (myininaya):

i did it on maple and it said it was infinity too

myininaya (myininaya):

that site said error

myininaya (myininaya):

lol cruffo's site is right

OpenStudy (anonymous):

will u chk out mine one too plz :

myininaya (myininaya):

i did but my calculator is also saying it diverges

OpenStudy (anonymous):

input at integral log(x)/sqrt(x) nd upper limit inf

myininaya (myininaya):

yes it says 4 but more evidence suggest it is infinity

OpenStudy (cruffo):

doesn't look like that calculator works with improper integrals, just definite ones. I tried it with your input, but that's not the correct answer.

myininaya (myininaya):

like maple and what we did above

OpenStudy (anonymous):

people this ques. is like headbanging .. I will b looking further into this lemme confirm what problem is with this problem

myininaya (myininaya):

this thread is starting to get pretty laggy

OpenStudy (cruffo):

In fact, I asked it to integrate x from 1 to inf, and it says the answer is approximatly -0.5

OpenStudy (anonymous):

ok.i got the ans

myininaya (myininaya):

from where?

myininaya (myininaya):

im still confident the answer is infinity (or diverges)

OpenStudy (anonymous):

working out.

OpenStudy (cruffo):

mathstina just wants us to shut up already :)

myininaya (myininaya):

lol

OpenStudy (anonymous):

ha

myininaya (myininaya):

omg i totally don't think that site works for improper integrals now try x^2, from 1 to inf

OpenStudy (anonymous):

watever, the ans is infinity right

myininaya (myininaya):

right lol